OpenStudy (anonymous):

if an object is dropped from a height of 85 feet the function h(t)=-16t^2+85 gives the height of the object after t seconds. approximately when will the object hit the ground? 85.00 69.00 0.33 2.30

OpenStudy (perl):

you solve h=0

OpenStudy (perl):

0 = -16t^2 + 85

OpenStudy (perl):

16t^2 = 85 t^2 = 85/16 t = + - sqrt( 85/16)

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

no C!?

OpenStudy (anonymous):

right?

OpenStudy (perl):

you dont own a calculator?

OpenStudy (anonymous):

on my phone...

OpenStudy (anonymous):

i got confused with the 85, but i honestly think its c

OpenStudy (perl):

C would be wrong

OpenStudy (perl):

read carefully what i wrote

OpenStudy (anonymous):

no idea......

OpenStudy (perl):

well im not supposed to give the answer. maybe someone else can help

OpenStudy (anonymous):

k

OpenStudy (perl):

sqrt( ) means take the square root

OpenStudy (anonymous):

yeah, i know that part its when ^2